OpenStudy (anonymous):

What is the equation in standard form that passes through the point (3, 4) and is parallel to the line y=-x+6

OpenStudy (anonymous):

Do you know that if a line is parallel to another line, then there slope is equal...

OpenStudy (anonymous):

Now tell me what is the slope of line : y = -x + 6??

OpenStudy (anonymous):

See, general equation we have is : y = mx + b Here, m is slope and b is y intercept...

OpenStudy (anonymous):

Compare the two equations and looks for green part and tell me the slope: \[y = \color{green}{m}x + b\] \[y = \color{green}{-1}x + 6\] What is the value of m when you compared?

OpenStudy (anonymous):

And one more thing, i will definitely need a reply from your side.. :)

OpenStudy (anonymous):

So is m in the first equation 1 if there parallel?

OpenStudy (anonymous):

What?? I did not get you. can you repeat?

OpenStudy (anonymous):

See, in this question you are given a line equation of which is : y = -x + 6.. But there is another line parallel to this line whose equation is not given which we have to find and this line is passing through points (3, 4) as given.. Getting the question??

OpenStudy (anonymous):